OpenStudy (anonymous):

how to solve heat equation using mathematica???does anybody know how to use programming???

OpenStudy (anonymous):

\[U_{t}=U _{xx} \] \[ 0<t<1\] \[0 \le t\]

OpenStudy (anonymous):

\[u(0,n)=u(1,n)=x\] \[0<n\] \[u(x,0)=x(1-x), 0 \le x \le 1\] \[\Delta n=0.1\space \Delta x=0.2\]